OpenStudy (anonymous):

A cyclist stops at a stop sign. After he checks the intersection, he begins pedaling again. He reaches a velocity of six meters per second in three seconds. What's the cyclist's acceleration? HELP ME PLEASE

OpenStudy (shreenithaa):

first write the formula

OpenStudy (shreenithaa):

what is it? i need to write it so i can talk u through this,okay?

OpenStudy (anonymous):

What do you mean the formula? Oh gosh I am not good at this!

OpenStudy (anonymous):

That is all that is there

OpenStudy (kl0723):

can you use velocity = intial velocity + accel x time? do you think you have all the data required?

OpenStudy (anonymous):

I have the initial velocity. That is 0. The velocity now is 2 m/s2 (That is squared. I do not know what direction the person is going) Time is 3 seconds I need the acceleration

OpenStudy (anonymous):

v=u+at

OpenStudy (kl0723):

since the cyclist stops and the starts to pedal again means that he starts from rest, then initial velocity is zero, you're given time and final velocity or max velocity in this case

OpenStudy (anonymous):

u=0 v=2m/s t=3 s a=?

OpenStudy (anonymous):

Yes?

OpenStudy (kl0723):

from the equation can you solve for Acceleration and then plug in your data?

OpenStudy (shreenithaa):

well, first of all, the formula is velocity equals distance divided by time acceleration is velocity final minus velocity intial over time final minus time intitial

OpenStudy (shreenithaa):

use this info to calculate ur math

OpenStudy (anonymous):

so it is 6 - 0 / and 3 - 0? so the acceleration is 2?

OpenStudy (shreenithaa):

velocity is 2 but u need final velocity and the starting velocity and final time and beginning time in order to get the acceleration we know that final time is 3 seconds

OpenStudy (anonymous):

I have the final velocity though. It is 2 m/s (Squared) the direction is not given. What I gave is what I have

OpenStudy (shreenithaa):

what if the starting time is 0 so now we need the initial velocity

OpenStudy (anonymous):

initial velocity is 0 m/s (Squared) since they are at a stop sign

OpenStudy (shreenithaa):

since the starting is 0 so the velocity must be 0 right?

OpenStudy (anonymous):

Yes

OpenStudy (shreenithaa):

so ur answer is acceleration is 2

OpenStudy (anonymous):

That is it?

OpenStudy (shreenithaa):

u were right the whole time sorry for the confusion

OpenStudy (anonymous):

The answer is just 2?

OpenStudy (shreenithaa):

no it has to have the units

OpenStudy (anonymous):

What are said units? Like are the from the problem, or?

OpenStudy (anonymous):

I just started physics recently, and boy is it hard!

OpenStudy (shreenithaa):

meters per second squared

OpenStudy (shreenithaa):

what grade r u in and state

OpenStudy (anonymous):

8th, texas

OpenStudy (anonymous):

Thank you, SO MUCH!

OpenStudy (shreenithaa):

8th California

OpenStudy (anonymous):

Cool! I'm just not very smart!

OpenStudy (anonymous):

2=0+a*3 a=2/3 m/s^2

OpenStudy (anonymous):

I er, have no idea what that means if we are being honest here

OpenStudy (shreenithaa):

wow same grad right!!! i moved to virginia where i had honors science so i learned this i was in cali for most of the time dont worry i am dumb in many places,but math and science are my specialty u'll find urs

OpenStudy (anonymous):

But thanks. Really!